solve for x
5=(x-2)/3+(x+1)/4

please show ur work so I can understand how to solve it myself. I'd really appreciate the help

Answer by mbmodak(32)

solve for x
5=(x-2)/3+(x+1)/4 here L.C.M. is 12 so we multiply both sides by 12 we get

5*12=((x-2)/3)*12+((x+1)/4)*12

60=(x-2)*4+(x+1)*3

60=4x-8+3x+3

60=4x+3x-8+3

60=7x-5

60+5=7x

65=7x

65/7=x ((65/7-2)/3+((65/7+1)/4)) = 5